# On three recent questions of Bourin and Lee on quadratic symmetric modulus and Euler-operator identity

### v1

@@ 4.3. A unitary-orbit Euler modulus inequality and singular value consequences
[…]
Proof.
[…]
@@
$4.$ Symmetric norms, antinorms, and Ky Fan consequences
@@
    For every symmetric antinorm $\|\cdot\|_{!}$ on $\mathbb{M}_{n}^{+}$ (in the sense of [3]),
@@
Proof.
Both inequalities follow from (1.3) and the (anti)triangle inequality on $\mathbb{M}_{n}^{+}$: a symmetric norm is subadditive, while a symmetric antinorm is superadditive. ∎
@@
Let $A,B,C\in\mathbb{M}_{n}$ and $m=1,\dots,n$. Denote by $\mu_{j}^{\downarrow}(\cdot)$ (resp. $\mu_{j}^{\uparrow}(\cdot)$) the singular values in nonincreasing (resp. nondecreasing) order. Then
@@
Proof.
Apply Corollary 4.12 to Ky Fan $m$-norms and Ky Fan $m$-antinorms on $\mathbb{M}_{n}^{+}$. ∎
@@ Acknowledgments
Teng Zhang is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No. xzy022024045).

### v2

@@ 8.1. A unitary-orbit bound for the Euler modulus
We now apply Lemma 8.1 with the concave function $f(t)=t^{1/2}$ to the Euler square-sum relation (1.6). This yields a three-term unitary-orbit domination of the Euler modulus by the pairwise moduli.
@@
As a first immediate consequence, taking any unitarily invariant norm in (8.1) and using the triangle inequality yields a norm inequality.
@@
We next record a singular-value refinement, obtained by combining Theorem 8.2 with the Weyl inequality for sums of positive semidefinite matrices.
@@
8.2. Symmetric norms, antinorms, and Ky Fan consequences
Finally, since Euler’s identity (1.6) is an exact equality of positive operators, it also yields further consequences for symmetric norms and antinorms (in the sense of [8]) on the positive semidefinite cone of $\mathbb{M}_{n}$, and in particular for Ky Fan sums of singular values.
@@ 8.2. Symmetric norms, antinorms, and Ky Fan consequences
    For every symmetric antinorm $\|\cdot\|_{!}$,
@@
Applying Corollary 8.5 to Ky Fan $m$-norms and Ky Fan $m$-antinorms gives the following eigenvalue-sum inequalities.
@@
Let $A,B,C\in\mathbb{M}_{n}$. Then for $m=1,\dots,n$,
@@ Acknowledgments
The author is very grateful to Professor J.-C. Bourin for reading this manuscript and for his valuable comments. This work is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No. xzy022024045).